# Worldline Formulations of Covariant Fracton Theories

### v1

@@
aDipartimento di Fisica e Astronomia “Augusto Righi”, Università di Bologna,
@@ 1 Introduction
The aim of this paper is to investigate possible worldline formulations of covariant fracton theories, continuing the broader program of reproducing the Batalin-Vilkovisky (BV) formulation of spacetime field theories [1, 2] from a first-quantized perspective.
@@
In recent years, exotic excitations with restricted mobility, called “fractons”, have been extensively studied, originally in the context of condensed matter physics [4, 5]. Fracton excitations appear in lattice spin models [6, 7], whose low-energy continuous limit [8] can be captured by effective field theories [9]. These models are characterized by single charged particles and dipoles with not only conserved charge, but also conserved dipole moment. The dipole-moment conservation implies for a single charged particle to be fixed in space, whereas dipoles are free to move.
@@
These features can be described by a pair of gauge fields, $\varphi$ and $A_{ij}$, where $i,j$ are space indices, with scalar gauge transformations $\delta\varphi=\partial_{t}\,\Lambda$ and $\delta A_{ij}=\partial_{i}\,\partial_{j}\,\Lambda$. Indeed, a coupling of the form $\varrho\,\varphi+J^{ij}\,A_{ij}$, for some fractonic matter current $\varrho$ and $J^{ij}$, leads to a higher-derivative continuity equation $\partial_{t}\,\varrho+\partial_{i}\,\partial_{j}\,J^{ij}=0$. It implies that both the charge $Q=\int\mathrm{d}^{D}x\,\varrho$ and the dipole moment $\vec{P}=\int\mathrm{d}^{D}x\,\vec{x}\,\varrho$ are conserved: $\frac{\mathrm{d}}{\mathrm{d}t}\,Q=0=\frac{\mathrm{d}}{\mathrm{d}t}\,\vec{P}$.
@@
These models break manifestly the Lorentz invariance. Nevertheless, they can be recovered in the framework of the gauge theory of the field $h_{\mu\nu}$, with gauge transformation (1.1), by setting $h_{0\mu}=\partial_{\mu}\,\varphi$, $h_{ij}=A_{ij}$ [10]. The covariant minimal coupling $j^{\mu\nu}\,h_{\mu\nu}$, for some covariant fractonic current $j^{\mu\nu}$, which, as a consequence of gauge invariance, satisfies the equation $\partial_{\mu}\,\partial_{\nu}\,j^{\mu\nu}=0$, becomes the original continuity equation, if $\varrho=\partial_{t}\,j^{00}+\partial_{i}\,j^{0i}$ and $J^{ij}=j^{ij}$. The label “fracton” for the covariant theory defined by the action (1.2) is then a bit misleading, but we maintain it, in accordance with existing literature on the topic ([3] and the subsequent works).
@@
Covariant fracton gauge theories cover a wider scope than the reasons why they were originally introduced. Indeed, on the one hand, they can be seen as a higher rank electrodynamics [10], because of the scalar gauge invariance (1.1) and the field strength (1.3), involving a single derivative. On the other hand, covariant fracton gauge theories can be thought of as a family of generalizations of linearized gravity [3, 11], where the invariance under linearized diffeomorphisms in the latter theory is restricted to longitudinal diffeomorphisms in the former. While the gauge invariance of linearized gravity reduces the propagating degrees of freedom to spin-2 polarizations only, in covariant fracton gauge theories propagating spin-1 and spin-0 excitations remain [12], which lead to instabilities [13]. They could be cured by considering suitable interactions – a first attempt is presented in [14]. The connection with gravity theories in the torsion formulation, as sketched in [15], may help clarify this issue.
@@
Further progress in understanding covariant fracton theories may benefit from exploring them through a different perspective in the context of worldline formalism [17, 18]. This approach offers an alternative to conventional second-quantized methods. On the one hand, it enables more efficient perturbative computations, as illustrated by recent calculations of heat-kernel coefficients and scattering amplitudes [19, 20, 21, 22]. On the other hand, being intrinsically a functional method, it is well-suited for non-perturbative analyses: in particular, it has proven useful in the study of strong-field phenomena, including the non-perturbative production of particle-antiparticle pairs [23, 24, 25, 26, 27].
Our goal is to construct first-quantized models capable of reproducing covariant fracton theories as spacetime theories. We will employ BRST quantization, which naturally yields the complete spectrum of spacetime fields required by the BV formalism, up to local redefinitions of the fields. Let us note that alternative worldline-based approaches to fractons have appeared in the literature [28, 29]. However, those works focus primarily on fracton matter, whereas our aim is to describe the covariant theory of the gauge field $h_{\mu\nu}$, defined by (1.3). Our manifestly Lorentz-invariant analysis set our framework apart, both conceptually and technically, from the earlier treatments.
@@
To describe more general gauge theories, the BRST formalism offers the most promising path forward. This technique actually finds its most successful application in the consistent coupling of worldline models to background fields, as illustrated in the seminal work [41]. There it was observed that a fully nilpotent BRST charge is necessary only on a restricted subspace of the extended Hilbert space. This relaxed condition allows for consistent interactions and has been applied successfully to couple massless and massive spin-2 particles to gravitational backgrounds [42, 43, 44] and massless/massive spin-1 particles to non-abelian/abelian backgrounds respectively [40, 24]. Relevant to this work, BRST methods have proven essential for the worldline formulation of a broader class of spacetime field theories. Notably, topological theories such as Chern-Simons theory have been formulated as worldline models [45], and even supersymmetric theories such as linearized $D=10$ super Yang-Mills and $D=11$ supergravity have been described using a first-quantized language [46, 47].
@@ 3.1 Tensor model
$\delta x^{\mu}$ $=\{x^{\mu},G\}=2\,\xi\,p^{\mu}+2\,\varepsilon\bar{\alpha}^{\mu\nu}\,p_{\nu}+2\,\bar{\varepsilon}\,\alpha^{\mu\nu}\,p_{\nu},$ (3.1.6)
$\delta p_{\mu}$ $=\{p_{\mu},G\}=0,$ (3.1.7)
$\delta\alpha^{\mu\nu}$ $=\{\alpha^{\mu\nu},G\}=i\,\varepsilon\,p^{\mu}\,p^{\nu}+i\,\phi\,\alpha^{\mu\nu},$ (3.1.8)
$\delta\bar{\alpha}^{\mu\nu}$ $=\{\bar{\alpha}^{\mu\nu},G\}=-i\,\bar{\varepsilon}\,p^{\mu}\,p^{\nu}-i\,\phi\,\bar{\alpha}^{\mu\nu},$ (3.1.9)
@@
$\delta e$ $=\dot{\xi}+i\,\bar{\varepsilon}\,u\,p^{2}-i\,\varepsilon\,\bar{u}\,p^{2},$ (3.1.10)
$\delta u$ $=\dot{\varepsilon}-i\,\varepsilon\,a+i\,\phi\,u,$ (3.1.11)
$\delta\bar{u}$ $=\dot{\bar{\varepsilon}}+i\,\bar{\varepsilon}\,a-i\,\phi\,\bar{u},$ (3.1.12)
$\delta a$ $=\dot{\phi}.$ (3.1.13)
@@
  • it must act on the (pre-quantization) phase-space variables, upon extending the Poisson brackets to account for the ghost variables, as the gauge transformations (3.1.6)–(3.1.9), with the ghost fields replacing the gauge parameters.
@@ 3.2 Vector model
$\delta x^{\mu}$ $=\{x^{\mu},G\}=\chi\,\alpha^{\mu}\,p_{\nu}\,\bar{\alpha}^{\nu}+\chi\,\bar{\alpha}^{\mu}\,p_{\nu}\,\alpha^{\nu}+2\,\varepsilon\,\bar{\alpha}^{\mu}\,p_{\nu}\,\bar{\alpha}^{\nu}+2\,\bar{\varepsilon}\,\alpha^{\mu}\,p_{\nu}\,\alpha^{\nu},$ (3.2.6)
$\delta p_{\mu}$ $=\{p_{\mu},G\}=0,$ (3.2.7)
$\delta\alpha^{\mu}$ $=\{\alpha^{\mu},G\}=i\,\chi\,p^{\mu}\,p_{\nu}\,\alpha^{\nu}+2\,i\,\varepsilon\,p^{\mu}\,p_{\nu}\,\bar{\alpha}^{\nu}+\tfrac{i}{2}\,\phi\,\alpha^{\mu},$ (3.2.8)
$\delta\bar{\alpha}^{\mu}$ $=\{\bar{\alpha}^{\mu},G\}=-i\,\chi\,p^{\mu}\,p_{\nu}\,\bar{\alpha}^{\nu}-2\,i\,\bar{\varepsilon}\,p^{\mu}\,p_{\nu}\,\alpha^{\nu}-\tfrac{i}{2}\,\phi\,\bar{\alpha}^{\mu},$ (3.2.9)
@@
$\delta\omega$ $=\dot{\chi}+4\,i\,H\,\bar{\varepsilon}\,u-4\,i\,H\,\varepsilon\,\bar{u},$ (3.2.10)
$\delta u$ $=\dot{\varepsilon}+2\,i\,H\,\chi\,u-2\,i\,H\,\varepsilon\,\omega-i\,\varepsilon\,a+i\,\phi\,u,$ (3.2.11)
$\delta\bar{u}$ $=\dot{\bar{\varepsilon}}-2\,i\,H\,\chi\,\bar{u}+2\,i\,H\,\bar{\varepsilon}\,\omega+i\,\bar{\,}\varepsilon a-i\,\phi\,\bar{u},$ (3.2.12)
$\delta a$ $=\dot{\phi}.$ (3.2.13)
@@
In a sense, the constraint $\ell$ plays a role in this model analogous to that of the Hamiltonian $H$ in model discussed in the previous section; this can also be inferred from the similarities between the gauge transformations (3.2.6)–(3.2.13) generated by $\ell$ and those of the previous model (3.1.6)–(3.1.13) generated by $H$ on the phase-space variables and on the gauge fields. [^12] One may wonder whether not-gauging the Hamiltonian constraint in a worldline model could lead to issues, as it certainly would in ordinary particle models. In the present case, we do not have a definite answer, and a deeper analysis, most likely involving the construction of the path integral, would be needed. Nonetheless, it would not be surprising to encounter instabilities already at the classical level, as this is a recurring issue in covariant fracton theories, even from a field-theoretic perspective [49, 13]. Choosing once again the normal ordering for the ghost sector, the nilpotent BRST charge is

### v2

@@
We develop worldline formulations of covariant fracton gauge theories. These are a one-parameter family of gauge theories of a rank-two symmetric tensor field, invariant under a scalar gauge transformation involving a double derivative. These theories, which can be interpreted as linearized gravity theories invariant under longitudinal diffeomorphisms, provide a covariant framework for studying Lorentz-breaking fracton quasiparticles, which are excitations with restricted mobility due to dipole-moment conservation.
We construct three worldline models. The first two are obtained by deducing their constraint structure directly from the spacetime gauge transformations. By applying BRST quantization, we show that these models reproduce the BV spectrum and the associated BRST transformations of two specific fracton theories. The third model is defined as a deformation of the second one: although free, it is analyzed by drawing inspiration from the standard treatment of interacting worldline systems, and is shown to capture almost the entire family of covariant fracton theories.
Finally, we discuss the gauge-fixing, comparing the BV-BRST spacetime perspective with the worldline analogue of the “Siegel gauge” employed in string field theory.
@@ 1 Introduction
The aim of this paper is to investigate possible worldline formulations of covariant fracton theories, continuing the broader program of reproducing the Batalin-Vilkovisky (BV) formulation of spacetime field theories [21, 22, 10] from a first-quantized perspective.
@@
In recent years, exotic excitations with restricted mobility, called “fractons”, have been extensively studied, originally in the context of condensed matter physics [68, 41, 64]. Fracton excitations appear in lattice spin models [55, 78], whose low-energy continuous limit [1] can be captured by effective field theories [67, 73]. These models are characterized by single charged particles and dipoles with not only conserved charge $Q=\int\mathrm{d}^{D-1}x\,\varrho$, but also conserved dipole moment $\vec{P}=\int\mathrm{d}^{D-1}x\,\vec{x}\,\varrho$, that is, $\frac{\mathrm{d}}{\mathrm{d}t}\,Q=0=\frac{\mathrm{d}}{\mathrm{d}t}\,\vec{P}$. Namely, the dipole-moment conservation implies for a single charged particle to be fixed in space, whereas dipoles are free to move.
These features are reproduced by a pair of gauge fields, $\varphi$ and $A_{ij}$, where $i,j$ are space indices, with gauge transformations
$\delta\varphi=\partial_{t}\,\Lambda,\quad\delta A_{ij}=\partial_{i}\,\partial_{j}\,\Lambda,$ (1.5)
@@
with $\Lambda$ a scalar parameter. Indeed, a coupling of the form $\varrho\,\varphi+J^{ij}\,A_{ij}$, for some fractonic matter current $\varrho$ and $J^{ij}$, leads to a higher-derivative continuity equation $\partial_{t}\,\varrho+\partial_{i}\,\partial_{j}\,J^{ij}=0$, which implies the conservation of both $Q$ and $\vec{P}$.
A concrete example of fractonic matter theory is studied in [69] (see also [67]), where the theory of a complex scalar field $\Phi$ is discussed. Such theory, besides the $U(1)$ invariance $\delta\Phi=i\,\alpha\,\Phi$, responsible of the charge conservation, enjoys also a dipole symmetry $\delta\Phi=i\,\beta_{i}\,x^{i}\,\Phi$, for some constant vector $\beta_{i}$, which implies the dipole-moment conservation. The gauging of such symmetries, as discussed in [33], requires the aforementioned pair of gauge fields $\varphi$ and $A_{ij}$, with gauge parameter $\Lambda=\alpha(x)+\beta_{i}(x)\,x^{i}$.
@@
Fractonic matter theories and fractonic gauge theories manifestly break the Lorentz invariance, since the gauge transformations (1.5) involve a different number of space and time derivatives. Nevertheless, the transformations (1.5) can be embedded in the covariant gauge transformation (1.1) of $h_{\mu\nu}$, by identifying $h_{0\mu}=\partial_{\mu}\,\varphi$, $h_{ij}=A_{ij}$ [32]. The covariant minimal coupling $j^{\mu\nu}\,h_{\mu\nu}$, for some covariant fractonic current $j^{\mu\nu}$, which, as a consequence of gauge invariance, satisfies the equation $\partial_{\mu}\,\partial_{\nu}\,j^{\mu\nu}=0$, becomes the original continuity equation, if $\varrho=\partial_{t}\,j^{00}+\partial_{i}\,j^{0i}$ and $J^{ij}=j^{ij}$. Therefore, the covariant theory provides a larger framework in which the genuine fracton models are embedded.
But since the distinctive feature of fractonic excitations – the reduced mobility – is by definition Lorentz-breaking, the label “fracton” for the covariant gauge theory defined by the action (1.2) is misleading. Nevertheless, we maintain it, in accordance with existing literature on the topic ([35] and the subsequent works). Moreover, we stress that by “covariant fracton gauge theories” we mean the pure Lorentz-invariant models of the gauge field $h_{\mu\nu}$, with action (1.2), without specifying any matter theory coupled to it. [^2] The Lorentz-breaking fractonic matter theory discussed in [69] can be covariantized in a natural way, as done in [3], thereby loosing the restriction in the mobility of the particle excitations, which is the distinctive feature of “true” fractons.
@@
Covariant fracton gauge theories are interesting per se, covering a wider scope than the reasons why they were originally introduced. Indeed, on the one hand, they can be seen as a higher rank electrodynamics [32], because of the scalar gauge invariance (1.1) and the field strength (1.3), involving a single derivative. On the other hand, covariant fracton gauge theories can be thought of as a family of generalizations of linearized gravity [35, 27], where the invariance under linearized diffeomorphisms in the latter theory is restricted to longitudinal diffeomorphisms in the former. While the gauge invariance of linearized gravity reduces the propagating degrees of freedom to spin-2 polarizations only, in covariant fracton gauge theories propagating spin-1 and spin-0 excitations remain [3], which lead to instabilities [2]. They could be cured by considering suitable interactions – a first attempt is presented in [31]. The connection with gravity theories in the torsion formulation, as sketched in [71], may help clarify this issue.
@@
Further progress in understanding covariant fracton theories may benefit from exploring them through a different perspective in the context of worldline formalism [72, 46, 14]. This approach offers an alternative to conventional second-quantized methods. On the one hand, it enables more efficient perturbative computations, as illustrated by recent calculations of heat-kernel coefficients and scattering amplitudes [18, 49, 6, 17]. On the other hand, being intrinsically a functional method, it is well-suited for non-perturbative analyses: in particular, it has proven useful in the study of strong-field phenomena, including the non-perturbative production of particle-antiparticle pairs [47, 20, 5, 61, 74].
Our goal is to construct first-quantized models capable of reproducing covariant fracton theories as spacetime theories. We will employ BRST quantization, which naturally yields the complete spectrum of spacetime fields required by the BV formalism, up to local redefinitions of the fields. Let us note that worldline-based approaches to (Lorentz-breaking) fractons models have appeared in the literature [42, 33]. Instead, our aim is to describe the covariant theory of the gauge field $h_{\mu\nu}$, defined by (1.3), from a wordline perspective, in a manifestly covariant way.
@@
To describe more general gauge theories, the BRST formalism offers the most promising path forward [58]. This technique actually finds its most successful application in the consistent coupling of worldline models to background fields, as illustrated in the seminal work [44]. There it was observed that a fully nilpotent BRST charge is necessary only on a restricted subspace of the extended Hilbert space. This relaxed condition allows for consistent interactions and has been applied successfully to couple massless and massive spin-2 particles to gravitational backgrounds [36, 37, 48] and massless/massive spin-1 particles to non-abelian/abelian backgrounds respectively [38, 20]. Relevant to this work, BRST methods have proven essential for the worldline formulation of a broader class of spacetime field theories. Notably, topological theories such as Chern-Simons theory have been formulated as worldline models [79], and even supersymmetric theories such as linearized $D=10$ super Yang-Mills and $D=11$ supergravity have been described using a first-quantized language [26, 25].
@@ 2 BV formulation of covariant fracton theories
The gauge theory described by the previous action is obviously invariant under a global Poincaré symmetry, since it is written in a manifestly covariant way, and it is also invariant under the following global scaling
$x^{\mu}\rightarrow\sigma^{-1}\,x^{\mu},\quad\partial_{\mu}\rightarrow\sigma\,\partial_{\mu},\quad h_{\mu\nu}\rightarrow\sigma^{\frac{D-2}{2}}\,h_{\mu\nu},$ (2.5)
such that any combination of the form $\mathrm{d}^{D}x\,\partial\,h\,\partial\,h$ is invariant. [^5] In [62] four models of fracton matter are considered in $2+1$ dimensions. They are obtained as the continuous limit of lattice models, involving fracton excitations due to dipole moment conservation. They are manifestly Lorentz-breaking since they involve a different number of spatial and temporal derivatives. These models are showed to enjoy subgroup of the conformal group as global symmetries.
@@ 3.1 Tensor model
$\ignorespaces\delta x^{\mu}$ $=\{x^{\mu},G\}=2\,\xi\,p^{\mu}+2\,\varepsilon\bar{\alpha}^{\mu\nu}\,p_{\nu}+2\,\bar{\varepsilon}\,\alpha^{\mu\nu}\,p_{\nu},$ (3.1.6a)
$\delta p_{\mu}$ $=\{p_{\mu},G\}=0,$ (3.1.6b)
$\delta\alpha^{\mu\nu}$ $=\{\alpha^{\mu\nu},G\}=i\,\varepsilon\,p^{\mu}\,p^{\nu}+i\,\phi\,\alpha^{\mu\nu},$ (3.1.6c)
$\delta\bar{\alpha}^{\mu\nu}$ $=\{\bar{\alpha}^{\mu\nu},G\}=-i\,\bar{\varepsilon}\,p^{\mu}\,p^{\nu}-i\,\phi\,\bar{\alpha}^{\mu\nu},$ (3.1.6d)
@@
$\ignorespaces\delta e$ $=\dot{\xi}+i\,\bar{\varepsilon}\,u\,p^{2}-i\,\varepsilon\,\bar{u}\,p^{2},$ (3.1.7a)
$\delta u$ $=\dot{\varepsilon}-i\,\varepsilon\,a+i\,\phi\,u,$ (3.1.7b)
$\delta\bar{u}$ $=\dot{\bar{\varepsilon}}+i\,\bar{\varepsilon}\,a-i\,\phi\,\bar{u},$ (3.1.7c)
$\delta a$ $=\dot{\phi}.$ (3.1.7d)
It is useful to briefly discuss the global symmetries of the worldline model. In particular, the model is invariant under the Poincaré group of the target space, which guarantees its relativistic invariance. If infinitesimal Lorentz transformations are parametrized by $\omega^{\mu}{}_{\nu}$, and spacetime translations by $k^{\mu}$, the corresponding transformations read [19]
$\delta x^{\mu}=\omega^{\mu}{}_{\nu}\,x^{\nu}+k^{\mu},\quad\delta p_{\mu}=\omega_{\mu}{}^{\nu}\,p_{\nu}.$ (3.1.8)
The rank-two symmetric tensor fields $\alpha^{\mu\nu}$ and $\bar{\alpha}_{\mu\nu}$ transform as
$\delta\alpha^{\mu\nu}=\omega^{\mu}{}_{\varrho}\,\alpha^{\varrho\nu}+\omega^{\nu}{}_{\varrho}\,\alpha^{\mu\varrho},\quad\delta\bar{\alpha}_{\mu\nu}=\omega_{\mu}{}^{\varrho}\,\bar{\alpha}_{\varrho\nu}+\omega_{\nu}{}^{\varrho}\,\bar{\alpha}_{\mu\varrho}.$ (3.1.9)
The worldline gauge fields are left invariant under both Lorentz transformations and translations. It then follows that the action (3.1.5) is invariant under the target-space Poincaré symmetry. We also note that the theory is manifestly invariant under the following global scale transformations:
$\tau\rightarrow\sigma^{p}\,\tau,\quad x^{\mu}\rightarrow\sigma^{-1}\,x^{\mu},\quad\alpha^{\mu\nu}\rightarrow\sigma^{m}\,\alpha^{\mu\nu},$ (3.1.10a)
$p_{\mu}\rightarrow\sigma\,p_{\mu},\quad\bar{\alpha}_{\mu\nu}\rightarrow\sigma^{-m}\,\bar{\alpha}_{\mu\nu},$ (3.1.10b)
$e\rightarrow\sigma^{-p-2}\,e,\quad u\rightarrow\sigma^{m-p-2}\,u,\quad\bar{u}\rightarrow\sigma^{-m-p-2}\,\bar{u},\quad a\rightarrow\sigma^{-p}\,a,$ (3.1.10c)
$H\rightarrow\sigma^{2}\,H,\quad\bar{L}\rightarrow\sigma^{-m-2}\,\bar{L},\quad L\rightarrow\sigma^{m+2}\,L,\quad J\rightarrow J,$ (3.1.10d)
for arbitrary numbers $m,p$.
A similar analysis applies to the worldline models discussed in the next sections and will not be repeated.
@@
  • it must act on the (pre-quantization) phase-space variables, upon extending the Poisson brackets to account for the ghost variables, as the gauge transformations (3.1.6a)–(3.1.6d), with the ghost fields replacing the gauge parameters.
@@ 3.2 Vector model
$\ignorespaces\delta x^{\mu}$ $=\{x^{\mu},G\}=\chi\,\alpha^{\mu}\,p_{\nu}\,\bar{\alpha}^{\nu}+\chi\,\bar{\alpha}^{\mu}\,p_{\nu}\,\alpha^{\nu}+2\,\varepsilon\,\bar{\alpha}^{\mu}\,p_{\nu}\,\bar{\alpha}^{\nu}+2\,\bar{\varepsilon}\,\alpha^{\mu}\,p_{\nu}\,\alpha^{\nu},$ (3.2.6a)
$\delta p_{\mu}$ $=\{p_{\mu},G\}=0,$ (3.2.6b)
$\delta\alpha^{\mu}$ $=\{\alpha^{\mu},G\}=i\,\chi\,p^{\mu}\,p_{\nu}\,\alpha^{\nu}+2\,i\,\varepsilon\,p^{\mu}\,p_{\nu}\,\bar{\alpha}^{\nu}+\tfrac{i}{2}\,\phi\,\alpha^{\mu},$ (3.2.6c)
$\delta\bar{\alpha}^{\mu}$ $=\{\bar{\alpha}^{\mu},G\}=-i\,\chi\,p^{\mu}\,p_{\nu}\,\bar{\alpha}^{\nu}-2\,i\,\bar{\varepsilon}\,p^{\mu}\,p_{\nu}\,\alpha^{\nu}-\tfrac{i}{2}\,\phi\,\bar{\alpha}^{\mu},$ (3.2.6d)
@@
$\ignorespaces\delta\omega$ $=\dot{\chi}+4\,i\,H\,\bar{\varepsilon}\,u-4\,i\,H\,\varepsilon\,\bar{u},$ (3.2.7a)
$\delta u$ $=\dot{\varepsilon}+2\,i\,H\,\chi\,u-2\,i\,H\,\varepsilon\,\omega-i\,\varepsilon\,a+i\,\phi\,u,$ (3.2.7b)
$\delta\bar{u}$ $=\dot{\bar{\varepsilon}}-2\,i\,H\,\chi\,\bar{u}+2\,i\,H\,\bar{\varepsilon}\,\omega+i\,\bar{\,}\varepsilon a-i\,\phi\,\bar{u},$ (3.2.7c)
$\delta a$ $=\dot{\phi}.$ (3.2.7d)
@@
In a sense, the constraint $\ell$ plays a role in this model analogous to that of the Hamiltonian $H$ in model discussed in the previous section; this can also be inferred from the similarities between the gauge transformations (3.2.6a)–(3.2.7d) generated by $\ell$ and those of the previous model (3.1.6a)–(3.1.7d) generated by $H$ on the phase-space variables and on the gauge fields. [^14] One may wonder whether not-gauging the Hamiltonian constraint in a worldline model could lead to issues, as it certainly would in ordinary particle models. In the present case, we do not have a definite answer, and a deeper analysis, most likely involving the construction of the path integral, would be needed. Nonetheless, it would not be surprising to encounter instabilities already at the classical level, as this is a recurring issue in covariant fracton theories, even from a field-theoretic perspective [70, 2]. Choosing once again the normal ordering for the ghost sector, the nilpotent BRST charge is